# The effect of friction on the dynamics of targeted energy transfer by symmetric vibro-impact dampers

### v1

@@ 4.4 Linear stability and bifurcation analysis
The numerical results shown in Figure (10) are consistent with the analytical findings. The computed solution branches of the $\mathcal{P}_{2}^{1}$ orbits are obtained through solving the nonlinear equations resulting from $\mathbf{G}^{\pm}\left(\mathbf{x}_{n},\mathbf{x}_{n+2}\right)=\mathbf{0}$ along with the periodicity condition (55) and (65). The stable branches are determined through an eigenvalue analysis of the numerically computed matrices $\mathbf{A}^{\pm}$. The computed solution branches present three types of bifurcations corresponding to the points where the stability property of each solution branch is lost. The behavior of the eigenvalues along these branches is illustrated in Figure (10), where the real and imaginary parts of the eigenvalues are displayed as functions of the parameter $\tilde{C}$. The most immediate observation from Figure (10), is that the evolution of the eigenvalues with increasing value of the amplitude $\tilde{C}$ generated from the numerical solutions of both equations $\mathbf{G}^{\pm}\left(\mathbf{x}_{n},\mathbf{x}_{n+2}\right)=\mathbf{0}$ coincide exactly as expected, since $\mathbf{G}^{+}$ and $\mathbf{G}^{-}$ describe the same motion in different coordinates. This provides an accurate prediction for the bifurcating behavior of the average solution $B$ from (65). Moreover, it is clear that $\tilde{C}_{\min}$ corresponds to an eigenvalue passing through $+1$, and it represents the point from which two solution branches emerge — only one of which is stable. As $\tilde{C}$ increases, the eigenvalues remain within the unit circle until $\tilde{C}_{\max}$ is reached. The stable branch between $\tilde{C}_{\min}$ and $\tilde{C}_{\max}$ corresponds to stable symmetric $\mathcal{P}_{2}^{1}$-orbits. At the critical value of $\tilde{C}_{\max}$, for which the corresponding eigenvalue passes through $+1$ again, the symmetry property is lost, and the stable asymmetric $\mathcal{P}_{2}^{1}$-orbits emerge from the SIM. The stability of the new branches is maintained as $\tilde{C}$ increases until a period-doubling bifurcation point is reached at $\tilde{C}=\tilde{C}_{\text{PD}}$ when an eigenvalue passes through $-1$.
@@ 5.1.1 Inactive absorber: Linear response
where the parameters $\Omega$ and $\tilde{g}$ are defined as in (8).
The critical value $C_{\text{cr}}$ depends not only on the normalized frequency $\Omega$ and the friction coefficient $\mu$, but also on the mass ratio $\epsilon$ and the cavity length $b$ (through $\tilde{g}$). The frequency response curve (FRC) in Figure (12) represents the maximum amplitude of the system’s response, i.e., $|q_{1}|_{\max}$, as a function of the frequency of the external excitation $f=\frac{\omega}{2\pi}$, for the linear oscillator with a sticking auxiliary mass. In this context, the critical displacement values of the main structure, $q_{1,\text{cr}}$, can be approximated according to (7) as
@@ 5.1.2 Purely sliding absorber: almost linear response
Another scenario involves the absorber sliding back and forth within the cavity without coming in contact with the side walls. This type of motion is directly related to whether the absorber’s velocity within the cavity satisfies its activation condition, as well as the initial conditions of the system, specifically where the flow starts in the slow variables space $\left(B,\,\tilde{C}\right)$. Although this case may initially seem similar to the frictionless scenario where $\mu=0$, it reveals one of the major effects of dry friction on the behavior of the system at hand. Specifically, during this motion, the auxiliary mass acts as a sliding absorber, with an additional energy dissipation due to friction. Moreover, the presence of friction generates an additional manifold within the slow variable space. The method of multiple scales can be employed again to predict the existence of equilibria on this manifold, following the same procedure as outlined previously.
@@
In other words, this case corresponds to setting $B=0$ in (20) and (25). According to (95), the time evolution of the normalized absolute velocity of the auxiliary mass during this particular type of motion, denoted by $\tilde{q}_{2}^{\prime}\approx w_{0}^{\prime}-v_{0}^{\prime}$, oscillates between ±$\pm h=\pm\frac{\pi}{2}\mu\tilde{g}$. Therefore, the expression for the relative displacement of the absorber within the cavity, $w_{0}$, is given by
@@
Here, the constant value of $B=0$ defines a new vertical manifold in the $(B,\tilde{C})$ parameter space. However, this manifold exists only within a limited amplitude range for $\tilde{C}$, as the sliding motion is constrained by the sidewalls of the cavity.
@@
$\begin{split}D_{1}\tilde{C}&=-\lambda\tilde{C}+\frac{16-\pi^{4}}{\pi^{4}}h\sqrt{\tilde{C}-h^{2}}-G\sqrt{\tilde{C}}\sin\left(\gamma\right)\,,\\ D_{1}\gamma&=-\frac{1}{2\tilde{C}}\left(G\sqrt{\tilde{C}}\cos\left(\gamma\right)-\frac{16-\pi^{4}}{\pi^{4}}h^{2}+2\sigma\tilde{C}\right)\,.\end{split}$ (104)
@@
[…]
@@ 5.1.3 Active vibro-impact absorber: Constant amplitude response
The influence of friction on the system’s response becomes clearer when comparing the frequency response curves of the frictionless case to those with friction. The main observation is that, in the constant amplitude response regime, friction reduces the active frequency range of the VI-NES and increases the activation threshold, i.e., $\tilde{C}_{\min}$, as shown in Figure (16). While these effects may appear unfavorable, they can be effectively controlled and managed during the design process, ensuring effective performance even when the ideal frictionless case cannot be achieved.
@@
Regarding the damping performance of the VI-NES with and without friction, each system offers superior performance in different frequency regions. The frictionless system performs better above the shifted resonance frequency, whereas the system with friction is more effective below it. As the frequency approaches resonance and the amplitude accordingly increases, the influence of friction diminishes, leading both systems to exhibit similar performance as the solution branches of both cases merge. At low amplitude levels, friction becomes non-negligible. As the amplitude approaches $q_{1,{\min}}$, i.e., $\tilde{C}\rightarrow\tilde{C}_{\min}$, the dissipation rate increases, thereby enhancing the overall damping effect of the VI-NES. This behavior is illustrated in Figure (17), where the free vibrations of the main structure are compared for both systems, shown in Figure (16). The results indicate that, although friction narrows the absorber’s activation frequency range and raises the activation threshold, it also enhances the decay rate and energy dissipation at lower response levels. This trade-off should be carefully considered in the design process to optimize the performance according to the system’s specific requirements.
@@ 5.2 Coexistence of solutions:
(c) Projection of the flow (green line) on the $(B,\tilde{C})$-plane. The red points correspond to the equilibria of the systems from (38) and (104). The black points represents the simulated steady-state amplitude.
(d) Projection of the flow (green line) on the $(\gamma,\tilde{C})$-plane. The blue line represents the flow of the system governed by (104) and the red point indicates the corresponding equilibrium.
@@
(c) Projection of the flow (green line) on the $(B,\tilde{C})$-plane. The red points correspond to the equilibria of the systems from (38) and (104). The black points represents the simulated steady-state amplitude.
(d) Projection of the flow (green line) on the $(\gamma,\tilde{C})$-plane. The blue line represents the flow of the system governed by (38) and the red point indicates the corresponding equilibrium.
@@ 5.3 Approximation of the frequency response curves
    nonlinear response with a purely sliding absorber (represented by the black lines), and
@@
Specifically, the brute force continuation method is used to trace the system’s solution for fixed parameters at a constant excitation level $E$, while varying the input frequency $f$. Identifying all possible solutions and their stability can be computationally expensive and highly sensitive to the step size $\Delta f$. While smaller steps improve accuracy, they also increase computational cost. For the presented results, the external excitation $e(t)$ is defined as a harmonic base excitation with a linearly and gradually changing frequency over the frequency range of interest. The system is then simulated for each value of the frequencies long enough for the trajectory to converge to a steady state, starting from an initial guess or previously converged trajectory. The simulated response level over the studied frequency range is then compared to the analytically determined branches. The results from forward and backward frequency sweeps, shown in Figures (22)-(23), confirm that the system’s response follows stable solution branches, indicating stable periodic regimes. Near resonance, although symmetric $\mathcal{P}_{2}^{1}$-orbits become unstable, the system transitions to other periodic solutions, such as asymmetric $\mathcal{P}_{2}^{1}$, $\mathcal{P}_{3}^{1}$, and $\mathcal{P}_{4}^{1}$-orbits, among others. Figures (22)-(23) include phase portraits of the absorber motion during these response regimes. This behavior aligns with the findings in Section 4, confirming a symmetry-breaking bifurcation followed by a period-doubling cascade around the unstable branch of the SIM of symmetric $\mathcal{P}_{2}^{1}$-orbits. Moreover, it demonstrates the existence of additional periodic motions beyond the period-doubling sequence of $\mathcal{P}_{2}^{1}$-orbits. As the amplitude increases, the solution progresses along stable branches interconnected through symmetry-breaking and period-doubling bifurcations, forming continuous segments of the frequency response curve. However, this continuity is not always maintained. Disruptions may occur due to the emergence of chaotic windows at the end of a period-doubling cascade or the occurrence of a fold bifurcation, where new periodic orbits appear. These phenomena can lead to gaps in the FRC, particularly near resonance, where multiple response regimes coexist. It should be noted that, in the frequency sweep results, the observed jumps between stable branches occur either due to a loss of stability at bifurcation points or due to the chosen frequency step size, with the latter requiring finer tuning to enhance the accuracy of the simulations. Moreover, the presence of stable solutions corresponding to purely sliding motion of the VI-NES eliminates the strongly modulated responses that typically arise just before VI-NES activation in the frictionless case (i.e., near the stable branches of periodic solutions with two impacts per period). In summary, the brute force continuation method has been effective in verifying the results, providing a reliable numerical approach for tracing solution paths and capturing qualitative changes in the system behavior.

### v2

@@ 4.4 Linear stability and bifurcation analysis
The numerical results shown in Figure (10) are consistent with the analytical findings. The computed solution branches of the $\mathcal{P}_{2}^{1}$ orbits are obtained through solving the nonlinear equations resulting from $\mathbf{G}^{\pm}\left(\mathbf{x}_{n},\mathbf{x}_{n+2}\right)=\mathbf{0}$ along with the periodicity condition (55) and (65). The stable branches are determined through an eigenvalue analysis of the numerically computed matrices $\mathbf{A}^{\pm}$. The computed solution branches present three types of bifurcations corresponding to the points where the stability property of each solution branch is lost. The behavior of the eigenvalues along these branches is illustrated in Figure (10), where the real and imaginary parts of the eigenvalues are displayed as functions of the parameter $\tilde{C}$. The most immediate observation from Figure (10), is that the evolution of the eigenvalues with increasing value of the amplitude $\tilde{C}$ generated from the numerical solutions of both equations $\mathbf{G}^{\pm}\left(\mathbf{x}_{n},\mathbf{x}_{n+2}\right)=\mathbf{0}$ coincide exactly as expected, since $\mathbf{G}^{+}$ and $\mathbf{G}^{-}$ describe the same motion in different coordinates. This provides an accurate prediction for the bifurcating behavior of the average solution $B$ from (65). Moreover, it is clear that $\tilde{C}_{\min}$ corresponds to an eigenvalue passing through +1+1, and it represents the point from which two solution branches emerge — only one of which is stable. As $\tilde{C}$ increases, the eigenvalues remain within the unit circle until $\tilde{C}_{\max}$ is reached. The stable branch between $\tilde{C}_{\min}$ and $\tilde{C}_{\max}$ corresponds to stable symmetric $\mathcal{P}_{2}^{1}$-orbits. At the critical value of $\tilde{C}_{\max}$, for which the corresponding eigenvalue passes through +1+1 again, the symmetry property is lost, and the stable asymmetric $\mathcal{P}_{2}^{1}$-orbits emerge from the SIM. The stability of the new branches is maintained as $\tilde{C}$ increases until a period-doubling bifurcation point is reached at $\tilde{C}=\tilde{C}_{\text{PD}}$ when an eigenvalue passes through 1-1.
@@ 5 Nonlinear response characteristics
5.1 Effect of friction on design-relevant performance metrics
This subsection presents closed-form expressions for the key design-relevant performance metrics of the VI-NES, derived from the Multiple Scales Method and the impact-map formulation, focusing on the symmetric $\mathcal{P}_{2}^{1}$-orbits. These metrics quantify how dry friction influences activation, stability, resonance-frequency shift, and decay. The expressions for the activation and amplitude bounds were derived in Sections 3-4, while the backbone curve and decay characteristics are obtained here by extending the procedure previously applied to the frictionless case in [33]. To maintain conciseness, intermediate algebraic steps are omitted and only the final closed-form expressions, reformulated in terms of the physical system parameters, and their physical implications are discussed.
Amplitude bounds and optimal activity range
The stability bounds of the symmetric $\mathcal{P}_{2}^{1}$-orbits define the main performance metrics of the VI-NES, corresponding to the amplitude level at which the absorber engages and the upper limit beyond which the optimal response for TET ceases. Reformulated in terms of the absolute displacement of the primary structure using the approximation
$q_{1}(t)\approx b\sqrt{\tilde{C}(\epsilon\omega_{0}t)}\sin(\omega_{0}t+\gamma)\,$, (92)
for a small mass ratio $\epsilon$, yields the following physical bounds:
$q_{1,\min}$ $=b\sqrt{\tilde{C}_{\min}}=\frac{\pi\,b\,h}{\sqrt{\pi^{2}+4R^{2}}}+\frac{2\,bR}{\sqrt{\pi^{2}+4R^{2}}}$
$=\mu\frac{\pi^{2}\ g}{2\omega_{0}^{2}\sqrt{\pi^{2}+4R^{2}}}+b\frac{2R}{\sqrt{\pi^{2}+4R^{2}}}\,$, (93)
$q_{1,\max}$ $=b\sqrt{\tilde{C}_{\max}}$
$=\frac{2\pi b}{\pi^{2}-4}\sqrt{\left(R\frac{\mu}{b}\frac{g}{\omega_{0}^{2}}\frac{\pi}{2}-\frac{2}{\pi}\right)^{2}+\left(R-\frac{\mu}{b}\frac{g}{\omega_{0}^{2}}\left(R^{2}+1-\left(\frac{\pi}{2}\right)^{2}\right)\right)^{2}}\,.$ (94)
Here, the use of $R$ preserves analytical clarity while highlighting the role of $r$ without unnecessary expansion. It should be noted that the tuning parameters might affect the normalized and physical amplitudes differently: for instance, increasing $b$ decreases $\tilde{C}_{\max}$ while increasing $q_{1,\max}$. This highlights the importance of basing design considerations on physical rather than normalized quantities. When defining the effective operating window of the symmetric $2$-IPP regime, however, both formulations lead to the same dimensionless measure,
$\xi=\frac{C_{\max}}{C_{\min}}=\sqrt{\frac{\tilde{C}_{\max}}{\tilde{C}_{\min}}}=\sqrt{\frac{\frac{q_{1,\max}^{2}}{b^{2}}}{\frac{q_{1,\min}^{2}}{b^{2}}}}=\frac{q_{1,\max}}{q_{1,\min}}\,$, (95)
which measures the width of the stable amplitude range and thus characterizes the energetic robustness of the VI-NES.
Decay rate and effective damping
Following [33], the free resonant response of the primary structure coupled to the VI-NES is examined. The amplitude evolution along the SIM is approximated using an upper linear bound of the slow dynamics,
$D_{1}\tilde{C}\leq-\tilde{\alpha}_{c}\tilde{C}+\tilde{K}_{c}\,$, (96)
where $\tilde{\alpha}_{c}$ and $\tilde{K}_{c}$ are defined in (39). Solving (96) yields the analytical envelope $\tilde{C}$, used to approximate the absolute displacement of the primary structure as
$q_{1,\text{app}}(t)=\sqrt{\left(\left(\frac{q_{10}}{\sin\gamma}\right)^{2}-\frac{\tilde{K}_{c}b^{2}}{\tilde{\alpha}_{c}}\right)e^{-\tilde{\alpha}_{c}\epsilon\omega_{0}t}+\frac{\tilde{K}_{c}b^{2}}{\tilde{\alpha}_{c}}}\sin\left(\omega_{0}t+\gamma\right)\,$, (97)
with $q_{10}:=b\sqrt{\tilde{C}(t=0)}\sin\gamma$}. Here, the decay rate $\tilde{\alpha}_{c}$ as well as the amplitude of the exponential term jointly characterize the overall dissipation strength. For the frictionless case studied in [33], the impact-induced damping ratio $D_{\text{NES}}\left(R,\,\epsilon\right)$, defined therein, provides a baseline dissipation level. In the presence of friction, the total dissipation increases as larger $\mu$ produces steeper decay and faster energy reduction. However, enhanced dissipation comes at the cost of a higher activation threshold, a trade-off that will be discussed in the following section.
Resonance shift and backbone curve
The nonlinear resonance properties of the coupled LO-VI-NES system are characterized through the backbone curve, derived by applying the phase-resonance condition within the slow-flow framework, following the procedure established in [33]. Substituting the condition $\gamma_{R}=-\frac{\pi}{2}+2j\pi$, $j\in\mathbb{Z}$, into the steady-state equations obtained by setting the right-hand side of (38) to zero yields the steady-state amplitude–forcing and detuning relations,
$G\sqrt{\tilde{C}_{R}}$ $=\tilde{\alpha}_{c}\tilde{C}_{R}+\tilde{\beta}_{c}\sqrt{\tilde{C}_{R}-\tilde{C}_{\text{min}}}-\tilde{K}_{c}\,$, (98)
$0$ $=\tilde{\alpha}_{\gamma}\tilde{C}_{R}+\tilde{\beta}_{\gamma}\sqrt{\tilde{C}_{R}-\tilde{C}_{\text{min}}}-\tilde{K}_{\gamma}\,$, (99)
with coefficients defined in (39). Here, the subscript $(\cdot)_{R}$ denotes quantities evaluated at resonance and should not be confused with the restitution-related parameter $R=\frac{1-r}{1+r}$. Rearranging (99) and reformulating in physical coordinates gives the resonance frequency
$\begin{split}f_{R}&=f_{0}(1+\epsilon\sigma_{R})\\ &=f_{0}\left(1+\frac{\epsilon}{2b\,q_{1,R}}\left(-\tilde{\beta}_{\gamma}\sqrt{q_{1,R}-q_{1,\min}}+\tilde{K}_{\gamma}\right)-\frac{\epsilon}{\frac{\pi^{2}}{4}+R^{2}}\right)\end{split}\,$, (100)
where the coefficients $\tilde{\beta}_{\gamma}$, $\tilde{K}_{\gamma}$, and $q_{1,\min}$ depend on $\mu$, $b$, and $R$ (see (39)). Equation (100) defines the backbone curve of the LO-VI-NES system, which connects the maxima of the frequency-response curves obtained for different excitation levels [48]. Here, a compact representation is preferred over a full expansion, as it retains parameter dependencies and preserves analytical transparency for design purposes.
Discussion
The analytical results provide quantitative insight into how dry friction affects the performance of the VI-NES within the optimal operating regime for targeted energy transfer. The influence of friction is neither inherently beneficial nor detrimental but depends on the excitation conditions and desired performance objectives. Increasing friction raises the activation threshold, leading to delayed engagement under harmonic excitation or earlier disengagement during free decay. This behavior is advantageous for harmonic excitation, ensuring VI-NES activation near resonance, but can limit energy dissipation during free decay, where a low activation threshold is preferable. Conversely, a low activation threshold can trigger early activation under harmonic loading, leading to higher-order periodic responses dominated by impacts and reduced frictional contribution near resonance.
While friction enhances dissipation, it simultaneously reduces the VI-NES activity bandwidth by narrowing both the frequency range and amplitude interval of efficient TET.
Overall, the influence of friction is strongly amplitude dependent: within the $2$-IPP regime, it introduces competing effects that lead to design trade-offs, whereas at higher amplitudes, its effect becomes negligible as the response converges toward the frictionless limit [46].
@@ 5.2.1 Inactive absorber: Linear response
where the parameters $\Omega$ and $\tilde{g}$ are defined as in (8). The critical value $C_{\text{cr}}$ depends not only on the normalized frequency $\Omega$ and the friction coefficient $\mu$, but also on the mass ratio $\epsilon$ and the cavity length $b$ (through $\tilde{g}$). The frequency response curve (FRC) in Figure (12) represents the maximum amplitude of the system’s response, i.e., $|q_{1}|_{\max}$, as a function of the frequency of the external excitation $f=\frac{\omega}{2\pi}$, for the linear oscillator with a sticking auxiliary mass. In this context, the critical displacement values of the main structure, $q_{1,\text{cr}}$, can be approximated according to (7) as
@@ 5.2.2 Purely sliding absorber: almost linear response
[…]
@@
In other words, this case corresponds to setting $B=0$ in (20) and (25). According to (104), the time evolution of the normalized absolute velocity of the auxiliary mass during this particular type of motion, denoted by $\tilde{q}_{2}^{\prime}\approx w_{0}^{\prime}-v_{0}^{\prime}$, oscillates between $\pm h=\pm\frac{\pi}{2}\mu\tilde{g}$. Therefore, the expression for the relative displacement of the absorber within the cavity, $w_{0}$, is given by
@@
[…]
@@
[…]
@@
To account for all possible sliding motions within the interval $\delta\in\left(0,\,1\right)$, the extreme cases, i.e., $\delta=0$ and $\delta=1$, are examined. These cases mark the critical transitions from sticking to sliding or from sliding to near-contact with the cavity walls, respectively. This analysis provides the amplitude range within which, the auxiliary mass behaves as a purely sliding absorber. By substituting the bounds of $\delta$ into (112), the existence range for this response regime is obtained as
@@ 5.2.3 Active vibro-impact absorber: Constant amplitude response
The influence of friction on the system’s response becomes clearer when comparing the frequency response curves of the frictionless case to those with friction. The main observation is that, in the constant amplitude response regime, friction reduces the active frequency range of the VI-NES and increases the activation threshold, i.e., $\tilde{C}_{\min}$, as shown in Figure (16). While these effects may appear unfavorable, they can be effectively controlled and managed during the design process, ensuring effective performance even when the ideal frictionless case cannot be achieved. Regarding the damping performance of the VI-NES with and without friction, each system offers superior performance in different frequency regions. The frictionless system performs better above the shifted resonance frequency, whereas the system with friction is more effective below it. As the frequency approaches resonance and the amplitude accordingly increases, the influence of friction diminishes, leading both systems to exhibit similar performance as the solution branches of both cases merge. At low amplitude levels, friction becomes non-negligible. As the amplitude approaches $q_{1,{\min}}$, i.e., $\tilde{C}\rightarrow\tilde{C}_{\min}$, the dissipation rate increases, thereby enhancing the overall damping effect of the VI-NES. This behavior is illustrated in Figure (17), where the free vibrations of the main structure are compared for both systems, shown in Figure (16). The results indicate that, although friction narrows the absorber’s activation frequency range and raises the activation threshold, it also enhances the decay rate and energy dissipation at lower response levels. This trade-off should be carefully considered in the design process to optimize the performance according to the system’s specific requirements.
@@ 5.3 Coexistence of solutions:
[…]
Figure 19: System response and phase portraits of the absorber’s periodic motion for $r=0.76$, $\mu=0.5$, $\tilde{g}=0.211$ and the external forcing parameters $G=2.47$ and $\sigma=-2.58$, simulated using (9)–(10). For the initial conditions $q_{1}(0)=q_{2}(0)=0\left[\text{m}\right]$ and $\dot{q}_{1}(0)=\dot{q}_{2}(0)=0\left[\text{m/s}\right]$, the absorber is purely sliding.
@@
[…]
Figure 20: System response and phase portraits of the absorber’s periodic motion for $r=0.76$, $\mu=0.5$, $\tilde{g}=0.211$ and the external forcing parameters $G=2.47$ and $\sigma=-2.58$, simulated using (9)–(10). For the initial conditions $q_{1}(0)=q_{2}(0)=0.02\left[\text{m}\right]$ and $\dot{q}_{1}(0)=\dot{q}_{2}(0)=0\left[\text{m/s}\right]$ , the auxiliary mass acts as an active vibro-impact absorber.
@@ 5.4 Approximation of the frequency response curves
    nonlinear response with a purely sliding absorber (represented by the black line), and
@@
Specifically, the brute force continuation method is used to trace the system’s solution for fixed parameters at a constant excitation level $E$, while varying the input frequency $f$. Identifying all possible solutions and their stability can be computationally expensive and highly sensitive to the step size $\Delta f$. While smaller steps improve accuracy, they also increase computational cost. For the presented results, the external excitation $e(t)$ is defined as a harmonic base excitation with a linearly and gradually changing frequency over the frequency range of interest. The system is then simulated for each value of the frequencies long enough for the trajectory to converge to a steady state, starting from an initial guess or previously converged trajectory. The simulated response level over the studied frequency range is then compared to the analytically determined branches.
@@
The results from forward and backward frequency sweeps, shown in Figures (22)-(23), confirm that the system’s response follows stable solution branches, indicating stable periodic regimes. Near resonance, although symmetric $\mathcal{P}_{2}^{1}$-orbits become unstable, the system transitions to other periodic solutions, such as asymmetric $\mathcal{P}_{2}^{1}$, $\mathcal{P}_{3}^{1}$, and $\mathcal{P}_{4}^{1}$-orbits, among others. Figures (22)-(23) include phase portraits of the absorber motion during these response regimes. This behavior aligns with the findings in Section 4, confirming a symmetry-breaking bifurcation followed by a period-doubling cascade around the unstable branch of the SIM of symmetric $\mathcal{P}_{2}^{1}$-orbits. Moreover, it demonstrates the existence of additional periodic motions beyond the period-doubling sequence of $\mathcal{P}_{2}^{1}$-orbits. As the amplitude increases, the solution progresses along stable branches interconnected through symmetry-breaking and period-doubling bifurcations, forming continuous segments of the frequency response curve. However, this continuity is not always maintained. Disruptions may occur due to the emergence of chaotic windows at the end of a period-doubling cascade or the occurrence of a fold bifurcation, where new periodic orbits appear. These phenomena can lead to gaps in the FRC, particularly near resonance, where multiple response regimes coexist. It should be noted that, in the frequency sweep results, the observed jumps between stable branches occur either due to a loss of stability at bifurcation points or due to the chosen frequency step size, with the latter requiring finer tuning to enhance the accuracy of the simulations. Moreover, the presence of stable solutions corresponding to purely sliding motion of the VI-NES eliminates the strongly modulated responses that typically arise just before VI-NES activation in the frictionless case (i.e., near the stable branches of periodic solutions with two impacts per period). In summary, the brute force continuation method has been effective in verifying the results, providing a reliable numerical approach for tracing solution paths and capturing qualitative changes in the system behavior.